OpenStudy (anonymous):

Help please Find the missing coefficient. (5d – 7)(5d – 6) = 25d2 + __ d + 42

OpenStudy (across):

Simply expand the LHS.

OpenStudy (bibby):

LHS being left hand side

OpenStudy (anonymous):

(4p _ 6)(4p + 6) a l6p2—36 b. Short Answer. 16

OpenStudy (anonymous):

a. 65 b. 5 c. - 5 d. - 65

OpenStudy (anonymous):

im going with B

OpenStudy (bibby):

You'd B wrong

OpenStudy (math&ing001):

Use FOIL on (5d – 7)(5d – 6), what do you get @MUNCH300 ?

OpenStudy (anonymous):

-5

OpenStudy (math&ing001):

Nope, show your work so I can help you.

OpenStudy (anonymous):

d. -65

OpenStudy (anonymous):

im doing it but im come how you get negative thoe i got 65 positive

OpenStudy (anonymous):

(5d – 7)(5d – 6)=5d(5d-6)-7(5d-6)=25d^2-30d-35d+42=25d^2-65d+42..done!!

OpenStudy (anonymous):

thanks